Information from the financial statements of Henderson-Niles Industries included the following at December 31, 2018:

Common shares outstanding throughout the year 100 million
Convertible preferred shares (convertible into 20 million shares of common) 70 million
Convertible 10% bonds (convertible into 19.0 million shares of common) $ 2,700 million


Henderson-Niles’s net income for the year ended December 31, 2018, is $940 million. The income tax rate is 40%. Henderson-Niles paid dividends of $2 per share on its preferred stock during 2018.

Required:
Compute basic and diluted earnings per share for the year ended December 31, 2018. (Enter your answers in millions (i.e., 10,000,000 should be entered as 10).)

Numerator / Denominator = Earnings per share
Basic $800 / 100 = $8.00
Diluted / = 0
Answer #1

Calculation of Basic Earning per share:

Net Income = 940 million

Preferred dividends = 70*2

= 140 per share

Basic EPS = (940-140)/100

= $8

Calculation of Diluted Earning per Share:

Diluted EPS =(940-140+140)/(100+20)

= $7.83